OpenStudy (anonymous):

divide using long division (5x^4-3x^3+x^2-6x+7) ÷ (x^2+1)

OpenStudy (jdoe0001):

\(\bf (5x^4-3x^3+x^2-6x+7) \div (x^2+1)\\ \quad \\\implies (5x^4-3x^3+x^2-6x+7) \div (x^2+0x+1)\\ \quad \\ \begin{array}{llll} \hline\\ 5x^4-3x^3+x^2-6x+7&| \end{array}\qquad x^2+0x-1\) what do you think would be our 1st number for the quotient?

OpenStudy (anonymous):

i'm not sure

OpenStudy (jdoe0001):

ok... what number multiplied by \(\bf x^2\) will give us \(\bf 5x^4\quad ?\)

OpenStudy (anonymous):

5^2

OpenStudy (jdoe0001):

25?

OpenStudy (anonymous):

i'm not sure i need alot of help haha

OpenStudy (anonymous):

can u explain it to me i'm a much better visual learner

OpenStudy (jdoe0001):

\(\bf a\cdot x^2=5x^4\implies a=\cfrac{5x^4}{x^2}\implies ?\) a value multiplied by \(\bf x^2\) to give us \(\bf 5x^4\).... so.. what do you think would be that value "a" ?

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=l6_ghhd7kwQ <--- this one covers it well

OpenStudy (anonymous):

so whats the answer?

OpenStudy (jdoe0001):

dunno